You drop a 2.50 kg textbook to a friend who stands on the ground at distance D = 10.0 m below. Your friend's outstretched hands are at distance d = 1.50 m above the ground.


Physics

(a) How much work Wg does the gravitational force do on the book as it drops to her hands?


(b) What is the change ΝU in the gravitational potential energy of the textbook-Earth system during the drop?


(c) If the gravitational potential energy of that system is zero at ground level, what is its potential energy U when the textbook is released?


(d) If the gravitational potential energy of that system is zero at ground level, what is its potential energy U when the textbook reaches the hands?


(e) How much work Wg is done on the textbook by its weight as it drops to your friend's hands if U is 100 J at the ground level.


(f) What is the change ΝU in the gravitational potential energy of the textbook-Earth system during the drop if U is 100 J at the ground level.


(g) Find U at the release point when U is 100 J at the ground level. Incorrect: Your answer is incorrect. J (h) Find U at the hands when U is 100 J at the ground level.
Answer #1

(a) W = mg*h = 2.5*9.8*(10-1.5) = 208.25 J


(b) Change = -W = -208.25 J


(c) P.E = 0+mg*h = 2.5*9.8*10 = 245 J


(d) P.E = 0 + mg*h = 2.5*9.8*1.5 = 36.75 J


(e) Work done remains the same = 208.25 J


(f) Change is also the same = -208.25 J


(g) P.E = 100 + mg*h = 345 J


(h) P.E = 100 + mg*h = 100 + 2.5*9.8*1.5 = 136.75 J